# Hybrid Near-field and Far-field Localization with Holographic MIMO

### v1

@@ VI-B1 Number of Cycles
Fig. 5 and Fig. 7 illustrate the angle estimation RMSE and the CPU time for different numbers of cycles. The RMSE is negatively correlated with the number of cycles, while the CPU time is positively correlated with the number of cycles, which demonstrates the trade-off relationship between localization accuracy and algorithm complexity.
@@ VI-B2 Sampling Spacing
Fig. 8 shows the average angle estimation RMSE of the proposed algorithm under different sampling spacing. Fig. 9 shows the CPU time of the proposed algorithm in three different models. It can be observed that as the angle sampling spacing reduces, the CPU time increases rapidly, while the estimation error first decreases and then remains fixed, which is consistent with previous analysis. Note that when the spacing is less than $\pi/40$, the accuracy does not improve, but the complexity still increases.
@@ VI-B3 RIS Size
We set the number of cycles as $C=50$ to ensure convergence. Fig. 1 and Fig. 11 show the range estimation RMSE and the CPU time of the proposed algorithm versus the number of RIS elements. As the number of RIS elements increases, the range estimation RMSE decreases, and CPU time increases. This is because the complexities of both the localization and optimization algorithms are positively correlated with the RIS element number. Also, a centimeter-level accuracy is attainable for a RIS with $900$ elements.
Figure 10: Range estimation performance of the NF users versus RIS size.11footnotemark: 1
Figure 11: CPU time of the proposed algorithm for different RIS sizes.
@@ Appendix A Proof of Proposition IV-A1
As mentioned in Section III-A, all previously received signals are used to estimate the locations, so we need to consider received signals in all previous $c$ cycles, as modeled in (31), to calculate the CRBs.
To obtain the expressions of CRBs, we first derive the expression of the Fisher information matrix (FIM) related to the unknown location parameters according to the received signals, denoted by $\bm{J}_{NF}$. The CRB, which gives the localization performance limit, is then derived from the FIM.

### v2

@@ VI-B1 Number of Cycles
Fig. 2 (a) illustrates the angle estimation RMSE for different numbers of cycles. It shows that for the hybrid and FF model, the estimation performance converges after $C=30$ cycles, while for the NF model, the result remains fixed after $C=35$ cycles, proving that the NF model is harder to converge. Fig. 2 (b) reveals the CPU time versus number of cycles for three models. The results show the proposed algorithm has the lowest complexity with FF model, followed by the hybrid model, and the highest complexity with NF model. This is because the FF model has the fewest candidate locations, while the NF model has the most.
Figure 7: (a) Angle estimation RMSE versus the number of cycles. (b) CPU time of the proposed algorithm versus number of cycles. [^2] Under the specified parameters at SNR = 5 dB, our method achieves an angle estimation error of $7^{\circ}$ and a range error of 0.2 m, with a computational latency of 0.15 s per update. This performance enables applications such as indoor navigation[44] and asset tracking[45].
Figure 8: (a) Localization of the NF user for different sampling spacing. (b) CPU time versus sampling spacing.
@@ VI-B2 Sampling Spacing
Fig. 8 (a) shows the average localization RMSE for NF users of the proposed algorithm under different sampling spacing. The localization RMSE is derived by computing user positions from estimated angles and ranges, then comparing them to the ground truth. Fig. 8 (b) shows the CPU time of the proposed algorithm in three different models. It can be observed that as the sampling spacing reduces, the CPU time increases rapidly, while the estimation error first decreases and then remains fixed, which is consistent with previous analysis. Note that when the spacing is less than $(0.25\text{m},\pi/40\ \text{rad})$, the accuracy does not improve, but the complexity still increases.
@@ VI-B3 RIS Size
We set the number of cycles as $C=50$ to ensure convergence. Fig. 9 (a) and Fig. 9 (b) show the range estimation RMSE and the CPU time of the proposed algorithm versus the number of RIS elements. As the number of RIS elements increases, the range estimation RMSE decreases, and CPU time increases. This is because the complexities of both the localization and optimization algorithms are positively correlated with the RIS element number. Also, a centimeter-level accuracy is attainable for a RIS with $900$ elements.
Figure 9: (a) Range estimation performance of the NF users versus RIS size. (b) CPU time of the proposed algorithm for different RIS sizes.
@@ Appendix A Proof of Proposition IV-A1
As mentioned in Section III-A, all previously received signals are used to estimate the locations, so we need to consider received signals in all previous $c$ cycles, as modeled in (31), to obtain the expressions of CRBs. We first derive the expression of the Fisher information matrix (FIM) related to the unknown location parameters according to the received signals, denoted by $\bm{J}_{NF}$. The CRB, which gives the localization performance limit, is then derived from the FIM.